HUBBLE SPACE TELESCOPE DAILY REPORT #1342 PERIOD COVERED: 0000Z (UTC) 03/07/95 - 0000Z (UTC) 03/08/95 Daily Status Report as of 067/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1 5413 (The Local ISM Flow Within the Heliosphere) The High Resolution Spectrograph was used to make observations to study the flow of neutral atoms in the interstellar medium by observing diffuse sky background in Lyman Alpha. A wavelength calibration image was also taken. Observations # 16, 18, and 19 were missing from a dump which was downloaded at 066/1200z (HSTAR #5150). The OSS reported that observation #15 ran long and an HRS-9 status buffer error message was received at 066/0357Z due to a carrousel out-of lock delay on observation #16. Observation #18 and 19 were lost due to a carrousel reset onboard delay. 1.2 Completed GHRS/1 5595 (GHRS ECH-A Wavelength Monitor: Cycle 4) The High Resolution Spectrograph was used to make wavelength spectral lamp calibration observations to measure the instrumental characteristics of the Echelle-A spectrograph. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.4 Completed WF/PC-2 5343 (WFC Imaging of Nearby Bright Quasars) The WF/PC-2 was used to make images of the bright quasar PG1402+261 and its host galaxy. The goal is to study the morphology of this galaxy, and to see how it differs from non-QSO galaxies. They also will be looking for the relationship between the properties of the host galaxy and the luminosity of the quasar. The observations were completed as planned, and no anomalies were reported. 1.5 Completed WF/PC-2 6140 (WF/PC-2 Cycle 4 Calibration: Ramp filters, Internal Calibration) The WF/PC-2 was used to make visual flat calibration images as part of a ramp filter calibration test. The observations were completed as planned and no problems were reported. 1.6 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6143 (WF/PC-2 Cycle 4 Photometric Calibration: Smaller Set, Part II) The WF/PC-2 was used to make single CCD calibration observations of standard star GRW+70D582 to monitor the photometric stability and quantum efficiency of the PC1 and WF3. The observations were executed as scheduled, and no anomalies were reported. 1.8 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of MARK699 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed FOS/BL 5581 (Pulsating White Dwarf DN Draconis) The FOS (Blue Side) was used in the rapid mode to cover the 17.1 hour beat period of the pulsating DA white dwarf DN Draconis (aka G226-29) to determine the index of the pulsational mode which will allow for the outer hydrogen layer to be determined. The observations completed nominally with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 5 Successful : 5 2.3 Significant HSTARs: None 2.4 Operations Notes: There are no significant operational problems to report. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-